Shift report 20030215 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Feb 15 2003 - 07:49:40 EST

  • Next message: Apache: "Shift report 20030215 08:00-16:00"
    ------------------------------------------------------------
    Date: Sat Feb 15 07:49:39 2003, Shift: 00:00-08:00
    Supervisor: MG, with: 
    Log book#14 pages 18 - 18
    
    Shift Summary:
    ------------------------------------------------------------
    1. Had a long store divided in three runs from 23:54 to 06:10. 3.5 hours storage in total. Gap cleaning started at this point.
    
    2. Finished MRS 90-degree with field 1050B. Finished FS 12-degree with field 1/8A. 
       Run 7900: MRS 90-degree,1050B
                 FS 12 degree,1/8 A
       Run 7901: MRS 90-degree,1050A
                 FS 12-degree,1/8 A
       Run 7902: MRS 90-degree,1050A
                 FS 12-degree,1/4 A
       Run 7903: Non TPC pedestal run.
       Run 7904: DVM pedestal run.
    
    3. Fixed problem with A1-yv anode in T3. I followed the same procedure in case of problems with A3-yv i.e. turned off A1 and F1, waited until currents went to zero, turned on first F1 then A1. No problem with other detectors.
    
    SITUATION: Keep current settings.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Feb 15 2003 - 07:50:18 EST